# Estimating effects of longitudinal modified treatment policies (LMTPs) on rates of change in health outcomes

### v1

@@ Abstract
Longitudinal data often contains time-varying outcomes measured at multiple visits and scientific interest may lie in quantifying the effect of an intervention on an outcome’s rate of change. For example, one may wish to study the progression (or trajectory) of a disease over time under different hypothetical interventions. We extend the longitudinal modified treatment policy (LMTP) methodology introduced in D´ıaz et al. (2023) to estimate effects of complex interventions on rates of change in an outcome over time. We exploit the theoretical properties of a nonparametric efficient influence function (EIF)-based estimator to introduce a novel inference framework that can be used to construct simultaneous confidence intervals for a variety of causal effects of interest and to formally test relevant global and local hypotheses about rates of change. We illustrate the utility of our framework in investigating whether a longitudinal shift intervention affects an outcome’s counterfactual trajectory, as compared with no intervention. We present results from a simulation study to illustrate the performance of our inference framework in a longitudinal setting with time-varying confounding and a continuous exposure.
@@ 1 Introduction
Longitudinal data often contains time-varying outcomes measured at each visit. Many methods for causal inference from longitudinal data aim to estimate effects on an outcome at the end of the study, effectively ignoring any additional outcome information that is available or treating intermediate outcomes as time-varying confounders. However, there are settings where the scientific interest lies not in an end-of-study effect but rather in the effect of an intervention on the outcome’s rate of change. For example, one may wish to investigate the effect of an intervention on the rate of progression of a disease, but the methods available to answer such questions are mostly limited to parametric models (e.g., linear models for the rate of change). In this work, we build on the flexible and nonparametric longitudinal modified treatment policy (LMTP) methodology to answer causal research questions that target effects on rates of change in outcomes over time.
@@
LMTPs quantify the effects of a broad class of interventions, including interventions that depend on the natural value of exposure (D´ıaz et al., 2023; Hoffman et al., 2024). The natural value refers to the value that a time-varying exposure would take at some time if the intervention was discontinued right before that time. LMTPs generalize “shift” interventions (e.g., where the natural value of exposure is shifted by a constant at every visit), “threshold” interventions (e.g., where the natural value of exposure is set to a threshold value when it fails to attain that threshold), and “stochastic” interventions (e.g., where the natural population- or individual-level values of exposure are modified) (Robins et al., 2004; Taubman et al., 2009; Muñoz and Van Der Laan, 2012; Haneuse and Rotnitzky, 2013; Young et al., 2014; D´ıaz and van der Laan, 2018; Kennedy, 2019; Sani et al., 2020). Static and dynamic regimes are also able to be studied as a special case. Since the LMTP framework accommodates such a wide range of interventions, it constitutes an ideal foundation to build upon. The LMTP approach has several other desirable qualities, particularly for longitudinal data. First, an LMTP approach allows researchers to evaluate policy-relevant interventions and encourages researchers to formulate estimands that correspond to interesting, complex research questions. Second, the LMTP methodology that is already developed properly accounts for time-varying confounding, a complication that is often encountered in longitudinal data. Third, the LMTP framework affords flexibility in the type of exposures being considered, allowing for continuous and categorical exposures as well as binary ones. Finally, the LMTP approach facilitates a focus on estimands that are more likely to satisfy the positivity assumption, which is key to reliable casual inference.
@@
In this paper, we extend the LMTP estimation methodology to estimate effects on rates of change in an outcome over time and develop a comprehensive framework to conduct relevant inference. The paper is organized as follows: Section 2 provides an introduction to LMTPs; Section 3 formalizes causal effects on rates of change; Section 4 proposes an inference framework and outlines its usefulness in investigating whether the rate of change in an outcome changes when an LMTP is implemented; Section 5 investigates the performance of the framework in a longitudinal setting using simulated data; and finally, Section 6 discusses limitations of the framework and proposes some future areas of work.
@@ 2 LMTPs
In this section, we provide an introduction to LMTPs and establish relevant notation. Our setup is similar to D´ıaz et al. (2023) but extended to accommodate a continuous time-varying outcome. Consider a sample of independent and identically distributed (iid) observations $Z_{1},\ldots,Z_{n}\sim P$, where distribution $P$ is contained in a nonparametric statistical model. Let $Z=(L_{1},A_{1},Y_{1},\ldots L_{\tau},A_{\tau},Y_{\tau})$. For discrete time $t\in\{1,\ldots,\tau\}$, $L_{t}$ is a vector of time-varying covariates, $A_{t}$ is a time-varying exposure, and $Y_{t}$ is a continuous time-varying outcome. For a random variable $X$, denote its history and future as $\overline{X}_{t}=(X_{1},\ldots,X_{t})$ and $\underline{X}_{t}=(X_{t},\ldots,X_{\tau})$, respectively. For notational simplicity, we use $\overline{X}$ and $\underline{X}$ to denote the complete history and future. Denote the history of all random variables until just before $A_{t}$ as $H_{t}=(\overline{A}_{t-1},\overline{L}_{t},\overline{Y}_{t-1})$.
@@
In this paper, we will often refer to this kind of shift intervention as our LMTP of interest; however, the theory developed applies to any LMTP that satisfies Definition 1. Other examples of LMTPs, including threshold interventions and LMTP stochastic interventions, are discussed in D´ıaz et al. (2023) and Hoffman et al. (2024).
@@ 3 Causal effects on rates of change
Theory for identification and nonparametric estimation of the vector of causal effects $\Delta$ follows straightforwardly from results established for any $\theta_{t}$ in D´ıaz et al. (2023). In brief, assumptions of positivity and strong sequential randomization are needed for identification based on a sequential regression formula. Two efficient influence function (EIF)-based estimators with many useful properties are proposed: the sequential doubly robust (SDR) estimator and the targeted minimum loss-based estimator (TMLE). Both of these estimators are $\sqrt{n}$-consistent and asymptotically normal. For these theoretical properties to hold, an additional assumption is needed: either $A_{t}$ must be discrete for all $t$ or if $A_{t}$ is continuous, $d(a_{t},h_{t})$ must be piecewise smooth invertible with respect to $a_{t}$. The framework that we propose in the next section relies heavily on the theoretical properties of these estimators. Nuisance functions in both the SDR estimator and TMLE are estimated using flexible data-adaptive techniques, such as Super Learner, that reduce the risk of model misspecification while still preserving proper convergence rates (Van der Laan et al., 2007). Additionally, these estimators are “doubly robust”, which refers to their ability to remain consistent even if some nuisance components are inconsistently estimated. Our implementation for estimating $\Delta$ follows straightforwardly from the functions available to estimate any $\theta_{t}$ in the lmtp R package (Williams and D´ıaz, 2023). In particular, we can estimate $\Delta$ by running the estimation algorithm for $\theta_{1},\dots,\theta_{\tau}$ (i.e., a loop of $\tau$ estimation runs) for both $d^{\prime}$ and $d^{\prime\prime}$.
@@ 4.2 Set up
Suppose we focus on constructing an estimator for $\theta$. For the purpose of illustrating this framework, we choose to use the SDR estimator instead of the TMLE because the SDR estimator is more robust to model misspecification in the sense that is is sequentially doubly robust, while the TMLE is only doubly robust (D´ıaz et al., 2023; Hoffman et al., 2024). Sequential double robustness guarantees that the estimator is consistent if either the outcome regression or treatment mechanism is consistently estimated for each time $t$. In contrast, double robustness only guarantees that the estimator is consistent if for some time $s$, all outcome regressions for $t>s$ and all treatment mechanisms for $t\leq s$ are consistently estimated. The only disadvantage of the SDR estimator is that unlike the TMLE, the SDR estimator can yield estimates of $\theta_{t}$ that fall outside the bounds of the parameter space. If this is an area of concern, one could consider using the TMLE instead. Even though the estimation algorithm for the TMLE differs from the SDR estimator, its construction based on the EIF allows a similar weak convergence result to be established so that the theory that we develop for inference still holds.
@@
where $g_{s}(a_{s}|h_{s})$ and $g_{s}^{d}(a_{s}|h_{s})$ are the densities of $A_{s}$ and $A_{s}^{d}$, respectively, conditional on $H_{s}=h_{s}$. Additional details on how the post-intervention density $g_{s}^{d}(a_{s}|h_{s})$ is defined can be found in D´ıaz et al. (2023).
The SDR estimator for any $\theta_{t}$ is constructed using a sequential regression algorithm (D´ıaz et al., 2023). This algorithm recursively regresses estimates of $\phi_{s+1,t}(Z;\underline{\eta}_{s})$ on $(A_{s},H_{s})$ for $s=t,\ldots,1$, starting with $\phi_{t+1,t}(Z;\underline{\eta}_{t+1})=Y_{t}$. The final output of this algorithm, $\phi_{1,t}(Z;\underline{\eta})$, is the (uncentered) EIF of $\theta_{t}$ in the nonparametric model. The SDR estimator for any $\theta_{t}$ under some intervention $d$ is then obtained by averaging the estimated EIF across all $i$ individuals, i.e.
@@
The SDR estimator $\widehat{\theta}_{n}$ satisfies a weak convergence result that follows directly from Theorem 4 in D´ıaz et al. (2023), which states that under proper convergence rates for the nuisance parameters, any $\widehat{\theta}_{t}$ weakly converges to a normal distribution with variance defined by the EIF.
@@
Assume conditions of Theorem 2 and 4 from D´ıaz et al. (2023) hold. Then,
@@ 5 Simulation study
We assessed the performance of our inference framework for investigating whether the natural outcome trajectory $\overline{\theta}^{\prime}$ differs from the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$ in terms of the rate of change. For each $n$, we repeated our simulation for a sequence of $\beta\in[0,1]$ to assess performance under both the null and alternative hypotheses. Figure 3 presents estimates of the natural outcome trajectory $\overline{\theta}^{\prime}$ and the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$, as well as the corresponding vector of causal effects $\Delta$, from a single simulated dataset when $\beta=1$ and $n=2500$. The simultaneous 95% confidence intervals of $\Delta$ capture the true non-zero causal effects, showing that our approach for estimation and simultaneous inference is performing as expected. Figure 4 presents the bias in the SDR estimator for $\Delta$ across several sample sizes $n$ when $\beta\in\{0,0.5,1\}$. As $n$ increases, bias converges to 0 as expected for each component of the SDR estimator of $\Delta$. Convergence is more rapid when $\beta$ is low and for components of the SDR estimator of $\Delta$ that are associated with earlier time points, suggesting that a larger sample size is needed for estimation using the Super Learner when there are stronger interaction effects and thus more variation in $Y_{t}$ in the underlying data generating process and when there are more time points to recurse through in the sequential regression algorithm used for estimation.
@@
Figure 3: Results from a single simulated dataset when $\beta=1$ and $n=2500$. Panel A visualizes estimates of the natural outcome trajectory $\overline{\theta}^{\prime}$ and counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$. Panel B visualizes an estimate of the difference trajectory (solid, black line) and the true difference trajectory (dashed, red line). Panel C visualizes an estimate of the vector of causal effects $\Delta$ and its simultaneous 95% confidence intervals (black), as well as the true vector of causal effects (red).

### v2

@@ Abstract
Longitudinal data often contains outcomes measured at multiple visits and scientific interest may lie in quantifying the effect of an intervention on an outcome’s rate of change. For example, one may wish to study the progression (or trajectory) of a disease over time under different hypothetical interventions. We extend the longitudinal modified treatment policy (LMTP) methodology introduced in Díaz et al. (2023) to estimate effects of complex interventions on rates of change in an outcome over time. We exploit the theoretical properties of a nonparametric efficient influence function (EIF)-based estimator to introduce a novel inference framework that can be used to construct simultaneous confidence intervals for a variety of causal effects of interest and to formally test relevant global and local hypotheses about rates of change. We demonstrate the utility of our framework in investigating whether a longitudinal shift intervention affects an outcome’s counterfactual trajectory, as compared with no intervention. We present results from a simulation study to illustrate the performance of our inference framework in a longitudinal setting with time-varying confounding and a continuous exposure. We also apply our inference framework to the Columbia Brain Health DataBank (CBDB) to examine the effect of shifting blood pressure on the progression of dementia.
@@ 1 Introduction
Longitudinal studies often measure outcomes repeatedly over time. Many methods for causal inference from longitudinal data aim to estimate effects on an outcome at the end of the study, effectively ignoring any additional outcome information that is available or treating intermediate outcomes as time-varying confounders. However, there are settings where the scientific interest lies not in an end-of-study effect but rather in the effect of an intervention on the outcome’s rate of change. For example, one may wish to investigate the effect of an intervention on the rate of progression of a disease, but the methods available to answer such questions are mostly limited to parametric models (e.g., linear models for the rate of change). In this work, we build on the flexible and nonparametric longitudinal modified treatment policy (LMTP) methodology to answer causal research questions that target effects on rates of change in outcomes over time.
@@
LMTPs quantify the effects of a broad class of interventions, including interventions that depend on the natural value of exposure (Díaz et al., 2023; Hoffman et al., 2024). The natural value refers to the value that a time-varying exposure would take at some time if the intervention was discontinued right before that time. LMTPs generalize “shift” interventions (e.g., where the natural value of exposure is shifted by a constant at every visit), “threshold” interventions (e.g., where the natural value of exposure is set to a threshold value when it fails to attain that threshold), and “stochastic” interventions (e.g., where the natural population- or individual-level values of exposure are modified) (Robins et al., 2004; Taubman et al., 2009; Muñoz and Van Der Laan, 2012; Haneuse and Rotnitzky, 2013; Young et al., 2014; Díaz and van der Laan, 2018; Kennedy, 2019; Sani et al., 2020). Static and dynamic regimes are also able to be studied as a special case. Since the LMTP framework accommodates such a wide range of interventions, it constitutes an ideal foundation to build upon. The LMTP approach has several other desirable qualities, particularly for longitudinal data. First, an LMTP approach allows researchers to evaluate policy-relevant interventions and encourages researchers to formulate estimands that correspond to interesting, complex research questions. Second, the LMTP methodology that is already developed properly accounts for time-varying confounding, a complication that is often encountered in longitudinal data. Third, the LMTP framework affords flexibility in the type of exposures being considered, allowing for continuous and categorical exposures as well as binary ones. Finally, the LMTP approach facilitates a focus on estimands that are more likely to satisfy the positivity assumption, which is key to reliable casual inference.
@@
Recent work on estimating LMTP effects has focused either on outcomes measured at the end of a study or survival outcomes (Díaz et al., 2023, 2024). In this paper, we extend the LMTP estimation methodology to estimate effects on rates of change in an outcome over time and develop a comprehensive framework to conduct relevant inference. The paper is organized as follows: Section 2 provides an introduction to LMTPs; Section 3 formalizes causal effects on rates of change; Section 4 proposes an inference framework and outlines its usefulness in investigating whether the rate of change in an outcome changes when an LMTP is implemented; Section 5 investigates the performance of the framework in a longitudinal setting using simulated data; Section 6 presents an illustrative application of the framework using the Columbia Brain Health DataBank (CBDB); and finally, Section 7 discusses limitations of the framework and proposes some future areas of work.
@@ 2 LMTPs
In this section, we provide an introduction to LMTPs and establish relevant notation. Our setup is similar to Díaz et al. (2023) but extended to accommodate a continuous outcome that is measured repeatedly over time. Consider a sample of independent and identically distributed (iid) observations $Z_{1},\ldots,Z_{n}\sim P$, where distribution $P$ is contained in a nonparametric statistical model. Let $Z=(L_{1},A_{1},Y_{1},\ldots L_{\tau},A_{\tau},Y_{\tau})$. For discrete time $t\in\{1,\ldots,\tau\}$, $L_{t}$ is a vector of time-varying covariates, $A_{t}$ is a time-varying exposure, and $Y_{t}$ is a continuous time-varying outcome. For a random variable $X$, denote its history and future as $\overline{X}_{t}=(X_{1},\ldots,X_{t})$ and $\underline{X}_{t}=(X_{t},\ldots,X_{\tau})$, respectively. For notational simplicity, we use $\overline{X}$ and $\underline{X}$ to denote the complete history and future. Denote the history of all random variables until just before $A_{t}$ as $H_{t}=(\overline{A}_{t-1},\overline{L}_{t},\overline{Y}_{t-1})$.
@@
In this paper, we will often refer to this kind of shift intervention as our LMTP of interest; however, the theory developed applies to any LMTP that satisfies Definition 1. Other examples of LMTPs, including threshold interventions and LMTP stochastic interventions, are discussed in Díaz et al. (2023) and Hoffman et al. (2024).
@@ 3 Causal effects on rates of change
Theory for identification and nonparametric estimation of the vector of causal effects $\Delta$ follows straightforwardly from results established for any $\theta_{t}$ in Díaz et al. (2023). In brief, assumptions of positivity and strong sequential randomization are needed for identification based on a sequential regression formula. Two efficient influence function (EIF)-based estimators with many useful properties are proposed: the sequential doubly robust (SDR) estimator and the targeted minimum loss-based estimator (TMLE). Both of these estimators are $\sqrt{n}$-consistent and asymptotically normal. For these theoretical properties to hold, an additional assumption is needed: either $A_{t}$ must be discrete for all $t$ or if $A_{t}$ is continuous, $d(a_{t},h_{t})$ must be piecewise smooth invertible with respect to $a_{t}$. The framework that we propose in the next section relies heavily on the theoretical properties of these estimators. Nuisance functions in both the SDR estimator and TMLE are estimated using flexible data-adaptive techniques, such as Super Learner, that reduce the risk of model misspecification while still preserving proper convergence rates (Van der Laan et al., 2007). Additionally, these estimators are “doubly robust”, which refers to their ability to remain consistent even if some nuisance components are inconsistently estimated. Our implementation for estimating $\Delta$ follows straightforwardly from the functions available to estimate any $\theta_{t}$ in the lmtp R package (Williams and Díaz, 2023). In particular, we can estimate $\Delta$ by running the estimation algorithm for $\theta_{1},\dots,\theta_{\tau}$ (i.e., a loop of $\tau$ estimation runs) for both $d^{\prime}$ and $d^{\prime\prime}$.
@@ 4.2 Set up
Suppose we focus on constructing an estimator for $\theta$. For the purpose of illustrating this framework, we choose to use the SDR estimator instead of the TMLE because the SDR estimator is more robust to model misspecification in the sense that is is sequentially doubly robust, while the TMLE is only doubly robust (Díaz et al., 2023; Hoffman et al., 2024). Sequential double robustness guarantees that the estimator is consistent if either the outcome regression or treatment mechanism is consistently estimated for each time $t$. In contrast, double robustness only guarantees that the estimator is consistent if for some time $s$, all outcome regressions for $t>s$ and all treatment mechanisms for $t\leq s$ are consistently estimated. The only disadvantage of the SDR estimator is that unlike the TMLE, the SDR estimator can yield estimates of $\theta_{t}$ that fall outside the bounds of the parameter space. If this is an area of concern, one could consider using the TMLE instead. Even though the estimation algorithm for the TMLE differs from the SDR estimator, its construction based on the EIF allows a similar weak convergence result to be established so that the theory that we develop for inference still holds.
@@
where $g_{s}(a_{s}|h_{s})$ and $g_{s}^{d}(a_{s}|h_{s})$ are the densities of $A_{s}$ and $A_{s}^{d}$, respectively, conditional on $H_{s}=h_{s}$. Additional details on how the post-intervention density $g_{s}^{d}(a_{s}|h_{s})$ is defined can be found in Díaz et al. (2023).
The SDR estimator for any $\theta_{t}$ is constructed using a sequential regression algorithm (Díaz et al., 2023). This algorithm recursively regresses estimates of $\phi_{s+1,t}(Z;\underline{\eta}_{s})$ on $(A_{s},H_{s})$ for $s=t,\ldots,1$, starting with $\phi_{t+1,t}(Z;\underline{\eta}_{t+1})=Y_{t}$. The final output of this algorithm, $\phi_{1,t}(Z;\underline{\eta})$, is the (uncentered) EIF of $\theta_{t}$ in the nonparametric model. The SDR estimator for any $\theta_{t}$ under some intervention $d$ is then obtained by averaging the estimated EIF across all $i$ individuals, i.e.
@@
The SDR estimator $\widehat{\theta}_{n}$ satisfies a weak convergence result that follows directly from Theorem 4 in Díaz et al. (2023), which states that under proper convergence rates for the nuisance parameters, any $\widehat{\theta}_{t}$ weakly converges to a normal distribution with variance defined by the EIF.
@@
Assume conditions of Theorem 2 and 4 from Díaz et al. (2023) hold. Then,
@@ 5 Simulation study
We assessed the performance of our inference framework for investigating whether the natural outcome trajectory $\overline{\theta}^{\prime}$ differs from the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$ in terms of the rate of change. For each $n$, we repeated our simulation for a sequence of $\beta\in[0,1]$ to assess performance under both the null and alternative hypotheses. Figure 3 presents estimates of the natural outcome trajectory $\overline{\theta}^{\prime}$ and the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$, as well as the corresponding vector of causal effects $\Delta$, from a single simulated dataset when $\beta=1$ and $n=2500$. The simultaneous 95% confidence intervals of $\Delta$ capture the true non-zero causal effects, while the global maximum test yields a significant result (test statistic of 27.5 with p-value <0.001<0.001), showing that our approach for estimation and inference is performing as expected. Figure 4 presents the bias in the SDR estimator for $\Delta$ across several sample sizes $n$ when $\beta\in\{0,0.5,1\}$. As $n$ increases, bias converges to 0 as expected for each component of the SDR estimator of $\Delta$. Convergence is more rapid when $\beta$ is low and for components of the SDR estimator of $\Delta$ that are associated with earlier time points, suggesting that a larger sample size is needed for estimation using the Super Learner when there are stronger interaction effects and thus more variation in $Y_{t}$ in the underlying data generating process and when there are more time points to recurse through in the sequential regression algorithm used for estimation.
@@
Figure 3: Results from a single simulated dataset when $\beta=1$ and $n=2500$. Panel A visualizes estimates of the natural outcome trajectory $\overline{\theta}^{\prime}$ and counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$. Panel B visualizes an estimate of the difference trajectory and the true difference trajectory. Panel C visualizes an estimate of the vector of causal effects $\Delta$ and its simultaneous 95% confidence intervals, as well as the true vector of causal effects.
@@
6 Illustrative application: the effects of shifted blood pressure on progression of dementia
We apply our proposed inference framework to investigate whether a longitudinal shift intervention on blood pressure affects the progression of dementia using the Columbia Brain Health DataBank (CBDB). The CBDB is an integrated data resource that combines Alzheimer’s disease (AD) research cohort data, electronic health records (EHR), neuroimaging, and biomarker information across the life course at Columbia University Irving Medical Center to understand the progression of Alzheimer’s disease and related dementias. 1,084 of the individuals in the CBDB are participants of the Columbia Alzheimer’s Disease Research Center (ADRC). Clinicians and research staff in the ADRC have conducted annual standardized cognitive evaluations on individuals with different etiologies of dementia, as well as those without a dementia diagnosis, since 2005 (Beekly et al., 2007). Our analysis focuses on 995 participants from this EHR-linked ADRC cohort.
Although cognitive evaluations are intended to be measured annually, the data is subject to complex patterns of missingness due to loss-to-follow up, drop-out/re-entry, refusal to respond, etc. As a result, there are steep drop offs in the number of participants available for analysis at each annual visit following the baseline visit. For example, from the 995 individuals with baseline visit 0, only 601 have visit 1, 513 have visit 2, and so on. In order to construct a sample over a 6-year time period that is sufficiently large enough for LMTP estimation, we binned the annual visits into 3 time-bins: baseline, 1-3 years since baseline, and 4-6 years since baseline. In cases with multiple visits in a bin, the latest visit was used. Additionally, since the current LMTP methodology only accommodates missingness from loss-to-follow up, we excluded all participants with any other patterns of missingness, leaving a total of 818 individuals in the final analysis. Extending our set up to accommodate loss-to-follow up follows straightforwardly from results established in Díaz et al. (2024). We now define the causal estimand $\theta_{t}$ as the expected counterfactual outcome at binned visit $t\in\{1,2,3\}$ under an intervention defined by function $d$ in a hypothetical world where there is no loss to follow-up. As a result, we must multiply the density ratio used in estimation by an indicator for being lost to follow-up, as well as the inverse probability of being lost to follow-up. The lmtp R package properly accommodates loss-to-follow up in the estimation of any $\theta_{t}$ by allowing users to specify indicators for being lost to follow-up, with 0 indicating being lost to follow-up at the next binned visit. The coarsening of observations into these bins is admittedly somewhat arbitrary and is not based on clinical considerations, so the results that we generate in this section should be interpreted only as illustrative of the proposed methodology.
In our analysis, the time-varying exposure of interest was systolic blood pressure. The outcome of interest was the Clinical Dementia Rating (CDR) Sum of Boxes score. Trained clinicians administer the CDR to patients to evaluate 6 cognitive domains (i.e., memory, orientation, judgment and problem solving, community affairs, home and hobbies, and personal care) (Hughes et al., 1982). The CDR yields both a global and Sum of Boxes score. The Sum of Boxes score ranges from 0 to 18 (with higher scores indicating worse performance) and is calculated by summing up the patient’s scores across the 6 domains. Studies have found that the Sum of Boxes score provides a finer gradation of impairment than the global score and is a promising primary endpoint for AD trials (Lynch et al., 2005; Coley et al., 2011). Studies have also demonstrated that the Sum of Boxes score can be used to stage dementia severity (O’Bryant et al., 2008, 2010). Finally, several studies have examined the rate of change in the Sum of Boxes score as a measure of the progression of dementia or AD (Rossetti et al., 2010; Tschanz et al., 2011; Williams et al., 2013). High systolic blood pressure in participants with a diagnosis of mild cognitive impairment (MCI) was associated with faster increases in the Sum of Boxes score (or in other words, faster progression of dementia) in the National Alzheimer’s Coordinating Center (NACC) Uniform Data Set (UDS) (Goldstein et al., 2013). In our analysis, body mass index (BMI) and time since baseline were time-varying covariates, while age at baseline, gender (categorized as male or female), and race/ethnicity (categorized as white non-Hispanic, Black non-Hispanic, Hispanic/Latino of any race, or other) were baseline variables. A descriptive summary of the exposure, outcome, and covariates in our study sample is provided in the Supplementary Materials (Table A1).
When a patient’s blood pressure is high, there are several possible interventions a physician may consider, including recommending lifestyle modifications and prescribing medications. We evaluate an additive shift LMTP that reduces the systolic blood pressure input by shift $\delta$ only if it is greater than or equal to threshold $\gamma$, or
$d(a_{t},h_{t})=\begin{cases}a_{t}-\delta,&\text{ if }a_{t}\geq\gamma\\ a_{t},&\text{ if }a_{t}<\gamma\end{cases}.$
Since physicians often recommend interventions when systolic blood pressure is above 130, the threshold for stage 1 hypertension, we present results when $\delta=10$ and $\gamma=130$ to investigate whether this LMTP affects the rate of change in the Sum of Boxes score, as compared with no intervention (Figure 7). Visual inspection suggests that the estimated counterfactual outcome trajectory increases less steeply than the natural outcome trajectory. Since the global maximum test yields a test statistic of 2.220 (p-value = 0.0497), we conclude that the two trajectories differ significantly in terms of their rates of change at the 0.05 significance level. Based on the maximum procedure for multiple testing, only the causal effect comparing 4-6 years later to baseline is significantly different from 0 (p-value = 0.0497). More specifically, the increase in the Sum of Boxes score from baseline to 4-6 years later is 0.309 (95% CI: 0.0003, 0.617) points lower, on average, under the LMTP versus no intervention. These results suggest that a shift intervention reducing systolic blood pressure may be protective against the progression of dementia, with benefits only appearing if enough time has passed.
Figure 7: Results from the illustrative application examining the effect of shifting blood pressure on the progression of dementia. The LMTP of interest reduces the systolic blood pressure input by shift $\delta=10$ only if it is greater than or equal to threshold $\gamma=130$. Panel A visualizes estimates of the natural outcome trajectory $\overline{\theta}^{\prime}$ and counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$. Panel B visualizes an estimate of the difference trajectory. Panel C visualizes an estimate of the vector of causal effects $\Delta$ and its simultaneous 95% confidence intervals.
Since one could envision a more aggressive intervention (i.e., with a greater shift or lower threshold), results for $\delta=20$ and $\gamma=130$, as well as for $\delta=10$ and $\gamma=120$, are also provided in the Supplementary Materials to demonstrate the impact of modifying the chosen LMTP on our conclusions (Figure A1). Decreasing the threshold to $\gamma=120$ has little impact on the estimated causal effects but yields a noticeable increase in their standard errors such that the global maximum test is no longer significant at the 0.05 significance level (test statistic of 1.669 with p-value = 0.160). Increasing the shift to $\delta=20$ leads to more negative estimates of the causal effects, especially comparing 4-6 years later to baseline, reflecting greater divergences between the natural and counterfactual outcome trajectories. It also yields even greater increases in the standard errors of the estimated causal effects. Once again, the global maximum test is no longer significant at the 0.05 significance level (test statistic of 2.070 with p-value = 0.067). These results highlight that even small modifications to the shift intervention under consideration may alter conclusions, particularly when the sample size is low.